OpenStudy (anonymous):

Last question of th enight, can anyone help me out? What is the center and radius of the following circle: (x – 5)2 + (y + 2)2 = 16

OpenStudy (anonymous):

Radius is 4 and the center is (5,-2)

OpenStudy (anonymous):

The equation for a circle is \[(x-x_0)^2 +(y - y_0)^2 = R^2 \] Where \[x_0\] is the x coordinate of the center, \[y_0\] is the y coordinate of the center, and \[R^2\] is the square of the radius of the circle.

OpenStudy (anonymous):

Now remember the Xo and the Yo always have a negative with them in the original equation so in this case, if the variable is positive it's the opposite.

OpenStudy (anonymous):

Hence, the y coordinate of the circle in this problem was"...(y+2)" tells me that y coordinate on the center of the circle is actually a negative 2

OpenStudy (anonymous):

So for my answer I have: (x−x0)^2+(y−y0)^2 = R^2 Radius = 4 Center = (5,-2) Is this right?

OpenStudy (anonymous):

Yep :) that's what I got! Oh p.s I helped you twice today. hahahha

OpenStudy (anonymous):

Lol thanks ma. I'll be sure to become your fan.

OpenStudy (anonymous):

No problem man. Now you can go rest ta-ta